OpenStudy (anonymous):

An automobile travels 34.o mpg of gasoline . how many kilometers does it travel per liter of gasoline ? use these equalities : 1 mile=1.61 kilometers ; 1 gallon = 3.79 liters

OpenStudy (ciarán95):

Let's work this out in steps: If the car does 34 mpg, that means 1 gallon of fuel will last for 34 miles of travel. So, 1 gallon = 34 miles. But, we know that 1 gallon = 3.79 liters.... So, now we can write 3.79 liters = 34 miles. If 34 miles can be covered in 3.79 liters, then 1 mile can be covered in 3.79/34 liters.... So, 1 mile = 3.79/34 liters, or approximately 0.111 liters if we work it out. However, we are told that 1 mile = 1.61 kilometers, so.... 1.61 kilometers = 0.111 liters. So, if 1.61 kilometers can be covered using 0.111 liters, then 1 kilometer can be covered by 0.111/1.61 liters, which you should be able to work out yourself. So, instead of mpg, this answer will be kilometers per liter of fuel. It seems complicated, but once you break it up into steps the maths is pretty simple. Hope that helps!

OpenStudy (anonymous):

i have been at this question forever could you just give the answer?

OpenStudy (ciarán95):

It's simply 0.111 divided by 1.61, it kind of defeats the purpose if I just give you the answer without understanding how to get to it yourself....it should be easy enough to work out using a calculator.

OpenStudy (anonymous):

i got 0.0689440994

OpenStudy (ciarán95):

Yes that's what I got as well....hopefully you can understand the steps now!

OpenStudy (anonymous):

your killing me lol how do i make it smaller

OpenStudy (ciarán95):

1 gallon = 34 miles But, 1 gallon = 3.79 liters, so 3.79 liters = 34 miles This means that 1 mile = 3.79/34 liters, or 0.111 liters (correct to 3 decimal places). But, 1 mile = 1.61 km, so 1.61 km = 0.111 liters SORRY, I MISREAD THE QUESTION!!! 0.0689440994 is the number of liters per kilometer, the number of kilometers covered per liter is 1.61/0.111. Hope that's OK now...and a bit simpler to follow! :)

OpenStudy (somy):

@backwoods15 i suppose its physics question?

OpenStudy (anonymous):

no chem

OpenStudy (somy):

oh interesting, looks like physics lol